# Vision Transformers in Precision Agriculture: A Comprehensive Survey

### v2

@@ 1 Introduction
Plant diseases pose a significant challenge to global agriculture, affecting crop yields, food security, and the economic well-being of farmers. Early and accurate detection is essential to effectively manage these diseases and minimize their impact [11]. Traditionally, diagnosing plant diseases has relied on labor-intensive manual inspections by experts or conventional machine learning techniques—both of which come with limitations. Manual inspection is not only time-consuming but also subject to human error and inconsistency, making it impractical for large-scale agricultural operations [48, 23]. To address these limitations, researchers are increasingly turning to automated solutions that harness recent advances in Machine Learning (ML) and computer vision. CNN, a widely used traditional Deep Learning (DL) method, have shown some success in automating the disease detection process [77, 43]. However, these models often struggle with scalability and with capturing complex spatial patterns in images—both of which are vital for accurate detection and classification of plant diseases across varied environmental conditions [50]. This has created a growing demand for more robust, scalable, and intelligent systems for plant disease detection. In response, several studies have explored the potential of DL to meet these challenges. Some existing surveys provide general overviews of DL approaches, highlighting their advantages over traditional methods and outlining the progress made in applying these techniques to plant disease detection tasks. For example, one study [73] reviews how DL and ML techniques have been employed for plant disease detection from image data. The study describes the transition from handcrafted features with the traditional DL methods to deep models like CNNs, DBNs, and segmentation models. It also discusses the implementation of object detection algorithms like YOLO and Faster R-CNN, and segmentation architectures like U-Net and Mask R-CNN, for pixel-level accuracy-based detection quality improvement. Another work [34] discusses the application of ML and DL techniques in the early detection and classification of plant diseases, highlighting their importance in improving agricultural yield and reducing crop loss. It presents an extensive list of ML techniques such as Naïve Bayes, k-Nearest Neighbors, Decision Trees, Support Vector Machines, Random Forests, and Multi-layer Perceptrons, and explores their applications in plant disease classification using environmental and image inputs. The study also discusses preprocessing techniques and image segmentation methods, demonstrating how ML has enabled the automation of disease diagnosis and soil testing from models trained with crop data. The contrast between classical ML and DL methods shows that DL outperforms ML in terms of accuracy and scalability. It emphasizes the importance of addressing current deficiencies and proposes DL as a promising direction for creating automatic, systematic, and scalable plant disease diagnostic systems that support both scientists and farmers in agricultural decision-making. Another overview of studies [2] explores the use of DL techniques for plant disease detection and management in agriculture. It addresses key challenges including dataset quality, imaging sensors, model generalizability, estimation of severity, and performance comparison with human capabilities. Commonly used datasets are categorized, along with their limitations in practical scenarios. The study outlines various DL techniques—particularly CNN-based models—used for image classification, object detection, and semantic segmentation. It identifies shortcomings in existing models’ ability to generalize to field conditions due to overfitting and imbalanced data. Additionally, significant research areas are highlighted, such as universal severity estimation, multi-disease detection, and model generalization across diverse conditions. A separate study [44] explores the application of DL in plant disease recognition and classification, emphasizing its critical role in agricultural productivity and food security. It illustrates how DL improves upon traditional image processing techniques through automated feature extraction and enhanced objectivity and efficiency. The paper discusses foundational concepts, model performance criteria, and the evolution of DL in plant disease recognition. It also addresses the need for data augmentation to manage limited datasets and the use of visualization methods to improve model interpretability. The study concludes with future challenges, such as the need for more varied datasets, enhanced model robustness, and the integration of hyperspectral imaging for early detection.
@@
The success of the Transformer architecture in NLP has revolutionized the field, establishing new benchmarks in tasks such as machine translation, text summarization, and language modeling [60]. By leveraging parallel processing and self-attention mechanisms, Transformers excel at capturing long-range dependencies and contextual relationships. This breakthrough has inspired researchers to explore the application of Transformers in other domains, including computer vision. It was soon recognized that the same properties that made Transformers effective in NLP—namely, their ability to model long-range dependencies and contextually relevant information—could also benefit image analysis tasks [19]. Unlike CNNs, which rely on inductive biases such as local connectivity and spatial hierarchies, Transformers offer a more flexible architecture with fewer assumptions about input structure. This flexibility enables them to learn rich, complex patterns directly from visual data, making them a promising alternative for various computer vision applications, including automated plant disease detection.
@@
The use of ViTs in agriculture is a novel and rapidly emerging research area. To the best of our knowledge, there are a few works whose focus is on ViTs; however, most of them mention it only briefly and do not provide a comprehensive analysis of their applications in precision agriculture. For example, a review [54] categorizes DL techniques applied to plant disease detection, organizing the literature into three tasks: classification, detection, and segmentation of plant diseases from leaf images. The review evaluates a broad range of DL models, with a strong focus on CNNs such as ResNet, VGG, and EfficientNet, alongside growing interest in ViTs. Another study [85] examines the use of DL and computer vision in plant disease detection, systematically analyzing the transition from traditional image processing to modern DL-based approaches. It assesses the suitability and performance of image processing techniques like thresholding, edge detection, and region-based segmentation in real-world agriculture. The study categorizes various DL architectures—including CNNs, Generative Adversarial Network (GAN), ViTs, and Vision Language Models—and highlights their scalability and practical implementation. It covers advanced models like YOLO and Faster R-CNN for real-time object detection, and discusses the emerging relevance of ViTs for capturing global context in image analysis. However, while ViTs are introduced, the review lacks a deep evaluation of specific architectures or performance benchmarks in agricultural contexts.
@@
  • A comprehensive survey is presented on the application of ViTs in the context of precision agriculture, systematically analyzing 42 high-impact peer-reviewed studies.
@@ 3.1 Definition of Inductive Biases
In supervised learning, when the data is frequently incomplete or noisy, inductive biases are essential for the model’s effectiveness. For example, in vision tasks, models such as CNNs assume that adjacent pixels are more likely to be correlated, which indicates an inductive bias based on the spatial structure of images. Similarly, in NLP, RNNs and Transformers utilize sequential and contextual relationships as biases to process and understand text. The type of inductive biases has a direct impact on a model’s performance on particular tasks. In computer vision—for example, in CNNs —some of the built-in inductive biases are locality and translation equivariance [18, 63, 22]. Locality assumes that important features, such as edges and textures, are restricted to small spatial regions within an image, while translation equivariance guarantees that these features are identified regardless of their position in the image. These biases enable CNNs to perform exceptionally well in tasks like image classification and object detection, where spatial hierarchies play a key role [39, 64].
@@ 3.2 Inductive Biases In CNNs and ViTs
CNNs have demonstrated remarkable effectiveness in modeling spatial hierarchies and capturing local dependencies within data. Their architecture, comprising convolutional layers succeeded by pooling layers, makes them very suitable for image recognition, object detection, and segmentation—areas where spatial relationships must be learned [71, 29]. The efficiency of CNNs is largely attributed to the strong inductive biases natively built into their design, including locality, two-dimensional neighborhood structure, and translation equivariance [19]. These are uniformly applied across all layers in the network, enabling CNNs to learn invariant and spatially-conscious feature representations.
@@
In particular, two dominant inductive biases—locality and weight sharing—serve as architectural constraints that significantly reduce the number of learnable parameters and enhance the model’s generalization capabilities [20]. Locality is the design principle whereby each neuron in a convolutional layer only connects to a small, spatially localized subset of the input—the receptive field. This local computation enables CNNs to effectively extract low-level features such as edges, corners, and textures. As data passes through deeper layers, local features are hierarchically combined to detect more abstract and global representations [25]. Weight sharing, meanwhile, ensures the same filters (kernels) are employed at different spatial locations, which makes the model equivariant to translation. As a result, inductively learned patterns in one region of an image are recognizable wherever they are, thereby making the model stronger and more efficient, significantly reducing the number of trainable parameters, enabling it to generalize well, and improving computational efficiency [16]. Together, these inductive biases not only reduce the model’s complexity but also add prior knowledge about the structure of visual data, which accelerates training and improves performance, especially in scenarios with limited labeled data. Although CNNs’ intrinsic inductive biases can be very effective when data is insufficient, they can be limiting when there is sufficient data. In this situation, we need to choose the most effective inductive biases—not necessarily all of them. [20]. The benefit of inductive biases tends to diminish with very large datasets, implying that transfer learning scenarios—where only a limited number of examples are available for the new distribution—provide a valuable context to assess the effectiveness and implementation of these biases [27]. It is due of these characteristics that a randomly initialized CNN can perform well in object localization tasks without any learning [12]. Despite all these advantages, the locality characteristic of CNNs causes them to degrade in their ability to capture long-range dependencies [20].
@@
ViTs have demonstrated remarkable effectiveness in modeling spatial hierarchies and capturing local dependencies within data. Unlike CNNs, which are specifically designed to operate with local data by spatially bounded filters, ViTs depend on self-attention mechanisms where every token (or image patch) can attend to any other token in the image regardless of their spatial positions [19]. In the ViT architecture, while Multi-Layer Perceptron (MLP) layers function locally and maintain a degree of translational equivariance, the self-attention layers are inherently global, enabling ViTs to integrate information across the entire spatial domain at every layer. Rather than relying on the strong inductive biases embedded within CNNs—such as locality, translation equivariance, and weight sharing—ViTs replace these constraints with global processing through multi-head self-attention. This design allows ViTs to learn flexible and context-dependent representations by dynamically attending to different parts of the image. As a result, ViTs are capable of incorporating broader contextual cues at earlier stages in the network, which can lead to superior performance in tasks that benefit from understanding long-range relationships between image regions. On the other hand, CNNs must build up such spatial relationships incrementally through repeated layers of local convolutions, making the learning process more architecture-biased than data-driven.
@@
However, this lack of built-in inductive biases also presents some challenges. The absence of explicit mechanisms for capturing local details can make ViTs less effective at modeling fine-grained patterns, especially in the early layers [61]. Additionally, ViTs exhibit high sensitivity to the choice of hyperparameters, such as the optimizer, learning rate schedule, and network depth, which makes their training more delicate and often requires large-scale datasets and computational resources [90]. Surprisingly, despite these architectural differences, recent results show that ViTs can implicitly learn some of the inductive biases that are commonly found in CNNs. For instance, early self-attention heads in ViTs prefer to attend to local regions, mimicking the effect of convolutional kernels [90]. This observation shows that with sufficient training data and good optimization, ViTs are capable of learning some of the beneficial priors that CNNs have by nature. While ViTs offer a more flexible and globally conscious architecture for visual representation learning, they lack the strong, human-crafted inductive biases that cause CNNs to excel so thoroughly in data-scarce regimes. This trade-off between flexibility and architectural bias highlights the importance of knowing when and how to use ViTs effectively—especially in terms of the scale and nature of the data in question.
@@ 3.3 Strategies to Mitigate the Lack of Inductive Biases in ViTs
Another strategy involves using knowledge distillation from a teacher model that possesses strong inductive biases. Distilling the knowledge in a neural network involves training a smaller neural network—known as the student or distilled model—to replicate the behavior of a larger, more complex model or a group of models, often referred to as the teacher or cumbersome model. The main goal is to transfer the broader understanding and generalization capability learned by the larger model into the smaller one [30]. In this approach, one or more pre-trained CNN models are often used as the teacher to transfer their inductive biases to the ViT student model [66]. This method has been shown to improve the performance of ViTs, particularly on smaller datasets.
Using hierarchical structures is another way to mitigate the lack of inductive biases by introducing a multi-stage architecture that processes visual information similarly to CNNs. Instead of attending globally from the start, these models apply self-attention within local windows and gradually merge or downsample tokens across layers, creating a spatial hierarchy [47]. This setup embeds locality and spatial coherence—key inductive biases of CNNs—allowing the model to first capture low-level features locally and then build up to more abstract, global representations [88]. Therefore, the most common approaches used to mitigate the lack of inductive biases in ViTs can be summarized as follows:
  • Convolutional Patch Embedding: It means replacing simple patch embedding with convolutional layers to preserve local structures in the image, or using recursive token aggregation—which behaves similarly to convolutions in capturing local spatial relationships—is a method employed in models like Tokens-to-Token ViT (T2T-ViT) [94].
@@ 4.1 Pure Models
A simple use-case of ViT is shown in the study by [52], which proposed a ViT-based approach for the automatic classification of strawberry diseases using transfer learning and fine-tuning techniques. The authors enhanced the ViT model by adding new layers such as ReLU activation, batch normalization, dropout, and a softmax classifier. Comparative evaluations against conventional CNN models like VGG16, VGG19, ResNet50V2, and MobileNet showed that the proposed ViT model, especially when combined with data augmentation, significantly outperformed other methods. In their best setting, the model achieved an accuracy and F1-score of 92.7% on the test set. The research highlights the strengths of Transformer-based architectures in precision agriculture, especially in complex visual tasks such as disease recognition from strawberry leaves, fruits, and flowers. The study also evaluated the proposed method on the PlantVillage dataset to assess generalizability, where it attained 98.9% accuracy, outperforming both the original ViT and previous state-of-the-art methods like Mask R-CNN and MCLCNN. The authors concluded that the ViT model’s ability to capture global image features makes it a powerful alternative to CNNs for agricultural disease detection tasks.
@@
As we discussed in Section 3, using hierarchical architecture is one of the possible solutions to mitigate the lack of inductive biases in ViTs. SwinT is a type of ViT designed to address some challenges associated with traditional ViTs, such as high computational cost for processing high-resolution images and the need for a hierarchical structure to effectively capture features at multiple scales [47]. There are some studies that used the SwinT model in precision agriculture. One use of SwinT was introduced by [98], in which the authors examined the use of Transformer models in accurate fruit ripeness classification. They addressed the problem of performing fruit selection and pickup effectively in the absence of professional labor. They created their own fruit datasets, including apples and pears, to train, test, and compare various Transformer models, like ViT, SwinT, and MLP-based models. The MLP model performed well for smaller models but was prone to overfitting when having higher capacity. The moving window-based self-attention in SwinT, as can be seen in Figure 5, facilitated better feature extraction of visual objects. The experiments demonstrated that SwinT yielded the best results, achieving a precision of 87.43% for fruit object detection. The work also combined the Transformer module and YOLO module to effectively classify the ripeness stages of apples and pears.
@@
[87] suggested an improved framework of the SwinT architecture to identify cucumber leaf diseases, to address complex backgrounds and the problem of limited sample sizes. The authors developed a SwinT-based and attention-guided GAN (STA-GAN) model, which is capable of producing diseased images without changing the background integrity. The leaf extraction module (Figure 6), which includes the proposed backbone network and Grad-CAM, was incorporated into the GAN to create STA-GAN (Figure 7). They employed a modified SwinT model, which uses an improved feature extraction technique of step-wise small patch embeddings and Coordinate Attention (CA), without adding any extra parameters. Their results showed that STA-GAN and the improved SwinT have the potential to minimize dependency on large sizes of manually annotated datasets, while at the same time improving disease detection.
@@
There are some studies whose main contribution is the use of new self-built or custom datasets. For instance, [8] presents a customized ViT model that was created to identify and categorize Java Plum leaf diseases. The dataset used in this experiment was self-collected, featuring six classes: Bacterial Spot, Brown Blight, Powdery Mildew, Sooty Mold, Dry, and Healthy leaves. The study showcases the capabilities of ViT for the improvement of plant disease detection technologies with respect to the example of Java Plum crops. The model that was developed in this work surpasses the performance of traditional CNN-based models and was able to provide better accuracy in detection. [17] explores the use of ViTs for plant disease detection using multispectral images taken in natural environments. This paper aims at assessing the performance of ViTs models in the area of plant disease recognition applied to multispectral imagery captured in natural conditions. The authors created a novel dataset named the “New Multispectral Dataset,” which contains pictures of five agricultural plants: avocados, tomatoes, lime, passionfruit, and gooseberries. This database includes pictures taken under both visible and near-infrared (NIR) capture modes, thus offering more information for parsing the diseases than the RGB-only capture. Four ViT models—ViT-S16, ViT-B16, ViT-L16, and ViT-B32—were trained from scratch and fine-tuned with ImageNet-21k weights. The research examined the roles of model size, patch size, and transfer learning on the ability of the models to classify diseases. The findings indicate that ViT’s model aimed at the B16 architecture and augmented with transfer learning provided the best results, with training and testing accuracies of 93.71% and 90.02%, respectively.
@@
[89] used the SwinT network to improve the identification of weeds in agriculture. To address the challenges posed by limited training data at scale, the authors employed a two-stage transfer learning strategy. First, they pre-trained the SwinT model on the ImageNet dataset, and then fine-tuned it on the Plant Seedlings dataset and a private dataset named MWFI (Maize/Weed Field Image). This strategy leveraged pre-trained network parameters to specialize in feature extraction for weed recognition. The proposed method achieved a recognition accuracy of 99.18%, outperforming other CNN-based architectures such as EfficientNetV2

### v3

@@ 1 Introduction
Plant diseases pose a significant challenge to global agriculture, affecting crop yields, food security, and the economic well-being of farmers. Early and accurate detection is essential to effectively manage these diseases and minimize their impact [11]. Traditionally, the diagnosis of plant diseases has depended on labor-intensive manual inspections by experts or on conventional Machine Learning (ML) techniques, both of which have inherent limitations.Manual inspection is not only time-consuming but also subject to human error and inconsistency, making it impractical for large-scale agricultural operations [57, 26]. Conventional ML methods are largely dependent on manual feature extraction, which limits their capacity to generalize across diverse environmental conditions, such as variations in lighting, background, and plant species. Moreover, these techniques are unable to effectively capture complex spatial patterns directly from raw image data. [24, 59]. To address these limitations, researchers are increasingly turning to automated solutions that harness recent advances in ML and computer vision. CNN, a widely used traditional Deep Learning (DL) method, have shown some success in automating the disease detection process [90, 52]. This has created a growing demand for more robust, scalable, and intelligent systems for plant disease detection. In response, several studies have explored the potential of DL to meet these challenges. Some existing surveys provide general overviews of DL approaches, highlighting their advantages over traditional methods and outlining the progress made in applying these techniques to plant disease detection tasks. For example, one study [85] reviews how DL and ML techniques have been employed for plant disease detection from image data. The study describes the transition from handcrafted features with the traditional DL methods to deep models like CNNs, DBNs, and segmentation models. It also discusses the implementation of object detection algorithms like YOLO and Faster R-CNN, and segmentation architectures like U-Net and Mask R-CNN, for pixel-level accuracy-based detection quality improvement. Another work [42] discusses the application of ML and DL techniques in the early detection and classification of plant diseases, highlighting their importance in improving agricultural yield and reducing crop loss. It presents an extensive list of ML techniques such as Naïve Bayes, k-Nearest Neighbors, Decision Trees, Support Vector Machines, Random Forests, and Multi-layer Perceptrons, and explores their applications in plant disease classification using environmental and image inputs. The study also discusses preprocessing techniques and image segmentation methods, demonstrating how ML has enabled the automation of disease diagnosis and soil testing from models trained with crop data. The contrast between classical ML and DL methods shows that DL outperforms ML in terms of accuracy and scalability. This study emphasizes the importance of addressing current deficiencies and proposes DL as a promising direction for creating automatic, systematic, and scalable plant disease diagnostic systems that support both scientists and farmers in agricultural decision-making. Another overview of studies [2] explores the use of DL techniques for plant disease detection and management in agriculture. It addresses key challenges including dataset quality, imaging sensors, model generalizability, estimation of severity, and performance comparison with human capabilities. Commonly used datasets are categorized, along with their limitations in practical scenarios. The study outlines various DL techniques, with a particular focus on CNN-based models, employed for image classification, object detection, and semantic segmentation. It highlights limitations in the ability of existing models to generalize to field conditions, primarily due to overfitting and imbalanced data. Additionally, significant research areas are highlighted, such as universal severity estimation, multi-disease detection, and model generalization across diverse conditions. A separate study [53] explores the application of DL in plant disease recognition and classification, emphasizing its critical role in agricultural productivity and food security. It illustrates how DL improves upon traditional image processing techniques through automated feature extraction and enhanced objectivity and efficiency. The paper discusses foundational concepts, model performance criteria, and the evolution of DL in plant disease recognition. It also addresses the need for data augmentation to manage limited datasets and the use of visualization methods to improve model interpretability. The study concludes with future challenges, such as the need for more varied datasets, enhanced model robustness, and the integration of hyperspectral imaging for early detection.
@@
Recent advancements in deep learning have demonstrated the potential of CNNs and their variants in accurately detecting and classifying plant diseases. However, despite their success, CNN-based models often struggle to capture global contextual information due to their inherent inductive biases and locality constraints. The introduction of the Transformer architecture has transformed the field of NLP, enabling significant advancements in machine translation, text summarization, and language modeling, and setting new performance benchmarks across a wide range of language understanding tasks [69]. By leveraging parallel processing and self-attention mechanisms, Transformers excel at capturing long-range dependencies and contextual relationships. This breakthrough has inspired researchers to explore the application of Transformers in other domains, including Computer Vision. It was soon recognized that the same properties that made Transformers effective in NLP, namely their ability to model long-range dependencies and contextually relevant information, could also be advantageous for image analysis tasks [21]. Unlike CNNs, which rely on inductive biases such as local connectivity and spatial hierarchies, Transformers offer a more flexible architecture with fewer assumptions about input structure. This flexibility enables them to learn rich, complex patterns directly from visual data, making them a promising alternative for various computer vision applications, including automated plant disease detection.
@@
The use of ViTs in agriculture is a novel and rapidly emerging research area. To the best of our knowledge, there are a few works whose focus is on ViTs; however, most of them mention it only briefly and do not provide a comprehensive analysis of their applications in precision agriculture. For example, a review [63] categorizes DL techniques applied to plant disease detection, organizing the literature into three tasks: classification, detection, and segmentation of plant diseases from leaf images. The review evaluates a broad range of DL models, with a strong focus on CNNs such as ResNet, VGG, and EfficientNet, alongside growing interest in ViTs. Another study [103] examines the use of DL and Computer Vision in plant disease detection, systematically analyzing the transition from traditional image processing to modern DL-based approaches. It assesses the suitability and performance of image processing techniques like thresholding, edge detection, and region-based segmentation in real-world agriculture. The study categorizes various DL architectures, such as CNNs, Generative Adversarial Network (GAN), ViTs, and Vision-Language Models, and highlights their scalability and potential for practical implementation. It covers advanced models like YOLO and Faster R-CNN for real-time object detection, and discusses the emerging relevance of ViTs for capturing global context in image analysis. However, while ViTs are introduced, the review lacks a deep evaluation of specific architectures or performance benchmarks in agricultural contexts.
@@
  • A comprehensive survey is presented on the application of ViTs in the context of precision agriculture, systematically analyzing 44 high-impact peer-reviewed studies.
@@ 3.1 Definition of Inductive Biases
In supervised learning, when the data is frequently incomplete or noisy, inductive biases are essential for the model’s effectiveness. In Computer Vision, for example in CNNs, some of the inherent inductive biases include locality and translation equivariance [20, 72, 25]. Locality assumes that important features, such as edges and textures, are restricted to small spatial regions within an image, while translation equivariance guarantees that these features are identified regardless of their position in the image. These biases enable CNNs to perform exceptionally well in tasks like image classification and object detection, where spatial hierarchies play a key role [47, 73]. Similarly, in NLP, RNNs and Transformers utilize sequential and contextual relationships as biases to process and understand text. The type of inductive biases has a direct impact on a model’s performance on particular tasks.
@@ 3.2 Inductive Biases In CNNs and ViTs
CNNs have demonstrated remarkable effectiveness in modeling spatial hierarchies and capturing local dependencies within data. Their architecture, which consists of convolutional layers followed by pooling layers, makes them well-suited for image recognition, object detection, and segmentation, where learning spatial relationships is essential [83, 33]. The efficiency of CNNs is largely attributed to the strong inductive biases natively built into their design, including locality, two-dimensional neighborhood structure, and translation equivariance [21]. These are uniformly applied across all layers in the network, enabling CNNs to learn invariant and spatially-conscious feature representations.
@@
In particular, two dominant inductive biases, locality and weight sharing, serve as architectural constraints that significantly reduce the number of learnable parameters and enhance the model’s generalization capabilities [22]. Locality is a design principle whereby each neuron in a convolutional layer connects only to a small, spatially localized subset of the input, known as the receptive field. This localized computation enables CNNs to effectively extract low-level features such as edges, corners, and textures. As data passes through deeper layers, local features are hierarchically combined to detect more abstract and global representations [28]. Weight sharing, meanwhile, ensures the same filters (kernels) are employed at different spatial locations, which makes the model equivariant to translation. As a result, inductively learned patterns in one region of an image are recognizable wherever they are, thereby making the model stronger and more efficient, significantly reducing the number of trainable parameters, enabling it to generalize well, and improving computational efficiency [17]. Together, these inductive biases not only reduce the model’s complexity but also add prior knowledge about the structure of visual data, which accelerates training and improves performance, especially in scenarios with limited labeled data. Although CNNs’ intrinsic inductive biases can be very effective when data is insufficient, they can be limiting when there is sufficient data. In this context, it is necessary to select the most effective inductive biases rather than employing all available ones [22]. The benefit of inductive biases tends to diminish with very large datasets, which implies that transfer learning scenarios, where only a limited number of examples are available for the new distribution, provide a valuable context for assessing the effectiveness and implementation of these biases [30]. It is due of these characteristics that a randomly initialized CNN can perform well in object localization tasks without any learning [13]. Despite all these advantages, the locality characteristic of CNNs causes them to degrade in their ability to capture long-range dependencies [22].
@@
ViTs have demonstrated remarkable effectiveness in modeling spatial hierarchies and capturing global dependencies within data. Unlike CNNs, which are specifically designed to operate with local data by spatially bounded filters, ViTs depend on self-attention mechanisms where every token (or image patch) can attend to any other token in the image regardless of their spatial positions [21]. In the ViT architecture, while Multi-Layer Perceptron (MLP) layers function locally and maintain a degree of translational equivariance, the self-attention layers are inherently global, enabling ViTs to integrate information across the entire spatial domain at every layer. Rather than relying on the strong inductive biases embedded within CNNs, such as locality, translation equivariance, and weight sharing, ViTs replace these constraints with global processing achieved through multi-head self-attention. This design allows ViTs to learn flexible and context-dependent representations by dynamically attending to different parts of the image. As a result, ViTs are capable of incorporating broader contextual cues at earlier stages in the network, which can lead to superior performance in tasks that benefit from understanding long-range relationships between image regions. On the other hand, CNNs must build up such spatial relationships incrementally through repeated layers of local convolutions, making the learning process more architecture-biased than data-driven.
@@
However, this lack of built-in inductive biases also presents some challenges. The absence of explicit mechanisms for capturing local details can make ViTs less effective at modeling fine-grained patterns, especially in the early layers [70]. Additionally, ViTs exhibit high sensitivity to the choice of hyperparameters, such as the optimizer, learning rate schedule, and network depth, which makes their training more delicate and often requires large-scale datasets and computational resources [108]. Surprisingly, despite these architectural differences, recent results show that ViTs can implicitly learn some of the inductive biases that are commonly found in CNNs. For instance, early self-attention heads in ViTs prefer to attend to local regions, mimicking the effect of convolutional kernels [108]. This observation shows that with sufficient training data and good optimization, ViTs are capable of learning some of the beneficial priors that CNNs have by nature. While ViTs offer a more flexible and globally conscious architecture for visual representation learning, they lack the strong, human-crafted inductive biases that cause CNNs to excel so thoroughly in data-scarce regimes. This trade-off between flexibility and architectural bias underscores the importance of understanding when and how to use ViTs effectively, particularly with respect to the scale and nature of the data involved.
@@ 3.3 Strategies to Mitigate the Lack of Inductive Biases in ViTs
Another strategy involves using knowledge distillation from a teacher model that possesses strong inductive biases. Distilling the knowledge in a neural network involves training a smaller neural network, known as the student or distilled model, to replicate the behavior of a larger, more complex model or ensemble of models, often referred to as the teacher or cumbersome model. The main goal is to transfer the broader understanding and generalization capability learned by the larger model into the smaller one [35]. In this approach, one or more pre-trained CNN models are often used as the teacher to transfer their inductive biases to the ViT student model [76]. This method has been shown to improve the performance of ViTs, particularly on smaller datasets.
Using hierarchical structures is another way to mitigate the lack of inductive biases by introducing a multi-stage architecture that processes visual information similarly to CNNs. Instead of attending globally from the start, these models apply self-attention within local windows and gradually merge or downsample tokens across layers, creating a spatial hierarchy [56]. This setup embeds locality and spatial coherence, which are key inductive biases of CNNs, allowing the model to initially capture low-level features locally and subsequently build more abstract, global representations [106]. Therefore, the most common approaches used to mitigate the lack of inductive biases in ViTs can be summarized as follows:
  • Convolutional Patch Embedding: Replacing simple patch embedding with convolutional layers to preserve local structures in the image, or employing recursive token aggregation, which functions similarly to convolutions in capturing local spatial relationships, is a method utilized in models such as Tokens-to-Token ViT (T2T-ViT) [113].
@@ 4.1 Pure Models
A simple use-case of ViT is shown in the study by [61], which proposed a ViT-based approach for the automatic classification of strawberry diseases using transfer learning and fine-tuning techniques. The authors enhanced the ViT model by adding new layers such as ReLU activation, batch normalization, dropout, and a softmax classifier. Comparative evaluations against conventional CNN models like VGG16, VGG19 [87], ResNet50V2 [34], and MobileNet [38] showed that the proposed ViT model, especially when combined with data augmentation, significantly outperformed other methods. In their best setting, the model achieved an accuracy and F1-score of 92.7% on the test set. The research highlights the strengths of Transformer-based architectures in precision agriculture, especially in complex visual tasks such as disease recognition from strawberry leaves, fruits, and flowers. The study also evaluated the proposed method on the PlantVillage dataset to assess generalizability, where it attained 98.9% accuracy, outperforming both the original ViT and previous state-of-the-art methods like Mask R-CNN [32] and MCLCNN. The authors concluded that the ViT model’s ability to capture global image features makes it a powerful alternative to CNNs for agricultural disease detection tasks.
@@
As we discussed in Section 3, using hierarchical architecture is one of the possible solutions to mitigate the lack of inductive biases in ViTs. SwinT is a type of ViT designed to address some challenges associated with traditional ViTs, such as high computational cost for processing high-resolution images and the need for a hierarchical structure to effectively capture features at multiple scales [56]. There are some studies that used the SwinT model in precision agriculture. One use of SwinT was introduced by [117], in which the authors examined the use of Transformer models in accurate fruit ripeness classification. They addressed the problem of performing fruit selection and pickup effectively in the absence of professional labor. They created their own fruit datasets, including apples and pears, to train, test, and compare various Transformer models, like ViT, SwinT, and MLP-based models. The MLP model performed well for smaller models but was prone to overfitting when having higher capacity. The moving window-based self-attention in SwinT, as can be seen in Figure 5, facilitated better feature extraction of visual objects. The experiments demonstrated that SwinT yielded the best results, achieving a precision of 87.43% for fruit object detection. This work also combined the Transformer module and YOLO [74] module to effectively classify the ripeness stages of apples and pears.
@@
[105] suggested an improved framework of the SwinT architecture to identify cucumber leaf diseases, to address complex backgrounds and the problem of limited sample sizes. The authors developed a SwinT-based and attention-guided GAN (STA-GAN) model, which is capable of producing diseased images without changing the background integrity. The leaf extraction module (Figure 6), which includes the proposed backbone network and Grad-CAM [81], was incorporated into the GAN to create STA-GAN (Figure 7). They employed a modified SwinT model, which uses an improved feature extraction technique of step-wise small patch embeddings and Coordinate Attention (CA), without adding any extra parameters. Their results showed that STA-GAN and the improved SwinT have the potential to minimize dependency on large sizes of manually annotated datasets, while at the same time improving disease detection.
@@
There are some studies whose main contribution is the use of new self-built or custom datasets. For instance, [8] presents a customized ViT model that was created to identify and categorize Java Plum leaf diseases. The dataset used in this experiment was self-collected, featuring six classes: Bacterial Spot, Brown Blight, Powdery Mildew, Sooty Mold, Dry, and Healthy leaves. The study showcases the capabilities of ViT for the improvement of plant disease detection technologies with respect to the example of Java Plum crops. The model that was developed in this work surpasses the performance of traditional CNN-based models and was able to provide better accuracy in detection. [19] explores the use of ViTs for plant disease detection using multispectral images taken in natural environments. This paper aims at assessing the performance of ViTs models in the area of plant disease recognition applied to multispectral imagery captured in natural conditions. The authors created a novel dataset named the “New Multispectral Dataset”, which contains pictures of five agricultural plants: avocados, tomatoes, lime, passionfruit, and gooseberries. This database includes pictures taken under both visible and near-infrared (NIR) capture modes, thus offering more information for parsing the diseases than the RGB-only capture. Four ViTs models, namely ViTs-S16, ViTs-B16, ViTs-L16, and ViTs-B32, were trained from scratch and subsequently fine-tuned using ImageNet-21k weights. The research examined the roles of model size, patch size, and transfer learning on the ability of the models to classify diseases. The findings indicate that ViTs’s model aimed at the B16 architecture and augmented with transfer learning provided the best results, with training and testing accuracies of 93.71% and 90.02%, respectively.
@@
[107] used the SwinT network to improve the identification of weeds in agriculture. To address the challenges posed by limited training data at scale, the authors employed a two-stage transfer learning strategy. First, they pre-trained the SwinT model on the ImageNet dataset, and then fine-tuned it on the Plant Seedlings dataset and a private dataset named MWFI (Maize/Weed Field Image). This strategy leveraged pre-trained network parameters to specialize in feature extraction for weed recognition. The proposed method achieved a recognition accuracy of 99.18%, outperforming other CNN-based architectures such as EfficientNetV2 [94].